OpenStudy (anonymous):

The range of a projectile depends not only on v0 and θ0 but also on the value g of the free-fall acceleration, which varies from place to place. In 1936, Jesse Owens established a world's running broad jump record of 8.09 m at the Olympic Games at Berlin (g = 9.8128 m/s2)) Assuming the same values of v0 and θ0, by how much would his record have differed ( in cm) if he had competed instead in 1956 at Melbourne (g = 9.7999 m/s2)?

OpenStudy (anonymous):

R1=u*u*(sin theta)/g1 R1=8.09 u*u*(sin theta)=R1*g1 =8.09*9.8128 R2=u*u*(sin theta)/g2 R2=8.09*9.8128/9.7999 =8.1006 difference=8.1006-8.0900 =0.0106 m = 1.06 cm